OpenStudy (anonymous):

whats 51 with a exponent of 1 ? whats 5 with a expontent of -2? whats 6 with a expontent of -3?

OpenStudy (anonymous):

\[b^{-n}=\frac{1}{b^n}\] and so \(5^{-2}=\frac{1}{5^2}=\frac{1}{25}\)